Entries and Schedules for Unfinished Jobs and Completed Jobs

Hildreth Company uses a job order cost system. The following data summarize the operations related to production for April, the first month of operations:

  1. Materials purchased on account, $2,690.
  2. Materials requisitioned and factory labor used:
    Job No. Materials Factory Labor
    101 $1,840 $2,910
    102 2,240 3,930
    103 1,490 1,920
    104 5,040 7,220
    105 3,200 5,500
    106 2,340 3,490
    For general factory use 630 4,310
  3. Factory overhead costs incurred on account, $3,510.
  4. Depreciation of machinery and equipment, $2,070.
  5. The factory overhead rate is $70 per machine hour. Machine hours used:
    Job No. Machine Hours
    101 25
    102 24
    103 24
    104 80
    105 19
    106 27
    Total 199
  6. Jobs completed: 101, 102, 103, and 105.
  7. Jobs were shipped and customers were billed as follows: Job 101, $7,800; Job 102, $9,420; Job 105, $14,760.

Required:

1. Journalize the entries to record the summarized operations. For a compound transaction, if an amount box does not require an entry, leave it blank.

Entries Description Debit Credit
a.
b.
c.
d.
e.
f.
g. Sale
g. Cost

2. Post the appropriate entries to T accounts for Work in Process and Finished Goods, using the identifying letters as transaction codes. Insert memo account balances as of the end of the month.

Work in Process
Bal.


Finished Goods
Bal.

3. Prepare a schedule of unfinished jobs to support the balance in the work in process account.

Hildreth Company
Schedule of Unfinished Jobs
Job Direct Materials Direct Labor Factory Overhead Total
$ $ $ $
Balance of Work in Process, April 30 $

4. Prepare a schedule of completed jobs on hand to support the balance in the finished goods account.

Hildreth Company
Schedule of Completed Jobs
Job Direct Materials Direct Labor Factory Overhead Total
$ $ $ $
Answer #1

1. Journalize the entries to record the summarized operations. For a compound transaction, if an amount box does not require an entry, leave it blank.

Entries Description Debit Credit
a. Raw materials inventory $2690
Accounts payable $2690
(To record materials purchased on account)
b. Work in process $41120
Factory overhead ($630+4310) $4940
Raw materials inventory $16780
Wages payable $29280
(To record materials and labor assigned)
c. Factory overhead $3510
Accounts payable $3510
(To record factory overhead costs incurred)
d. Factory overhead $2070
Accumulated depreciation- Machinery and equipment $2070
(To record depreciation expense)
e. Work in process (199*$70) $13930
Factory overhead $13930
(To record overhead applied)
f. Finished goods $29470
Work in process $29470
(To record jobs transferred to finished goods)
g. Sale Accounts receivable ($7800+9420+14760) $31980
Sales $31980
(To record sales)
g. Cost Cost of goods sold $24380
Finished goods sold $24380
(To record cost of goods sold)

b. Work in process= $1840+2240+1490+5040+3200+2340+2910+3930+1920+7220+5500+3490= $41120

Raw materials inventory= $1840+2240+1490+5040+3200+2340+630= $16780

Wages payable= $2910+3930+1920+7220+5500+3490+4310= $29280

f. Finished goods= Job 101+102+103+105

Job 101= $1840+2910+(25*$70)= $6500

102= $2240+3930+(24*$70)= $7850

103= $1490+1920+(24*$70)= $5090

105= $3200+5500+(19*$70)= $10030

Finished goods= $6500+7850+5090+10030= $29470

g. Cost of goods sold= Job 101+102+105

= $6500+7850+10030= $

2. Post the appropriate entries to T accounts for Work in Process and Finished Goods, using the identifying letters as transaction codes. Insert memo account balances as of the end of the month.

Work in Process
b. $41120 f. 29470
e. 13930
Bal. $25580
Finished Goods
f. 29470 g. 24380
Bal. 5090

3. Prepare a schedule of unfinished jobs to support the balance in the work in process account.

Hildreth Company
Schedule of Unfinished Jobs
Job Direct Materials Direct Labor Factory Overhead Total
104 $5040 $7220 (80*$70)= 5600 $17860
106 2340 3490 (27*$70)= 1890 7720
Balance of Work in Process, April 30 $25580

4. Prepare a schedule of completed jobs on hand to support the balance in the finished goods account.

Hildreth Company
Schedule of Completed Jobs
Job Direct Materials Direct Labor Factory Overhead Total
103 $1490 $1920 (24*$70)= $1680 $5090